aliciasara90:

1. A commercial prawn fisherman recorded the number of prawns he took from each trap. The average number of prawns per trap was 230 with a standard deviation of 22. What number of prawns per trap would you expect in the interval symmetrical about the mean where 80% of the numbers would be found?

mhanifa:

The numbers would be between 10% and 90% mark, do you know how to find z-scores?

mhanifa:

The equation: \[z = (x - \mu)/\sigma \]

aliciasara90:

idk how to find the z scores i know they are -1.28 and 1.28

aliciasara90:

but idk how to get them

mhanifa:

-1.28 is incorrect, you get them from z- tables http://www.z-table.com/

aliciasara90:

how do i know which score it is from the table

mhanifa:

10% is less than mean, so to the left, hence is negative 90% is positive for same reason

mhanifa:

You have \[\mu = 230, \sigma = 22, z - two different scenarios\] You need to find out x, there will be two of them, one for minimum and one for maximum mark

mhanifa:

Can you find x based on what you have?

aliciasara90:

how did u get -2.33 and 1.28

aliciasara90:

im not sure how to read thetable

mhanifa:

See the attached, here I highlighted the one for 10%. By the way you are correct, they are -1.28 and 1.28

1 attachment
mhanifa:

The closest number to 10% or 0,1 is 0,1003

mhanifa:

See if you get it

aliciasara90:

oh okay i undestand now that part now thank you

aliciasara90:

so then z=x-m/o

aliciasara90:

to find x?

mhanifa:

Yes, get two values of x

mhanifa:

1.28 = (x - 230)/22 -1.28 = (x - 230)/22 Solve each for x

mhanifa:

What did you get?

mhanifa:

Any help?

aliciasara90:

thanks i got 201 to 258

mhanifa:

Yes, I believe it is correct

aliciasara90:

thanks for your help

mhanifa:

No problem